Give proof to show:

Start with xp X and Iw and Tw to prove Y P
Answer #1

part 1:

l=r * p

==>dl/dt= r * (dp/dt)

time rate of change in momentum=force

hence dl/dt=r*F

but r*F=torque

hence proved.

part 2:

l=I*w

==>dl/dt=I*(dw/dt)

but dw/dt=time rate of change of angular velocity=angular acceleration

hence

dl/dt=moment of inertia*angular acceleration=torque applied

hence proved.